An elevator is moving upward at a constant speed of 5.55 m/s when a bolt suddenly breaks free from the bottom of the elevator. If the elevator was 91.8 meters above the bottom of the elevator shaft when the bolt breaks loose, it will take how many seconds to reach the bottom of the elevator shaft.
1
Expert's answer
2016-09-20T10:01:15-0400
t = S/v = 91.8/5.55 = 16.54 s